# Dichotomy theorem separating complete integrability and non-integrability of isotropic spin chains

### v3

@@ Abstract
Here, in the latter case we exclude all $k$-local conserved quantities with $13\leq k\leq N/2-$ 1. Our result establishes the dichotomy of local conserved quantities (all or nothing), excluding a partially-integrable system YCS24-1 ; YCS24-2 with a finite number of nontrivial local conserved quantities. In addition, the criterion of this dichotomy is whether a single quantity is zero or not, which is easy to calculate. This shows a high contrast to the Grabowski-Mathieu conjecture GM95-2 ; GP21 and its partial proof by Hokkyo Hok25 , where we need to solve search problems on operator equations to confirm non-integrability.
@@ I.1 Yang-Baxter equation and the quantum inverse scattering method
$[T(\lambda),T(\lambda^{\prime})]$ $=0,$ (S.4)
@@ I.3 Boost operator
where we used $R(0)=\Pi$ and $\left.\frac{d}{d\mu}R(\mu)\right|_{\mu=0}=\Pi h$. This relation is sometimes called the Sutherland equation. We reassign subscripts as $1\to a$, $2\to i$, $13\to i+$ 1, and multiply $\prod_{j\leq i-1}R_{ja}(\lambda)$ from left and $\prod_{j\geq i+2}R_{ja}(\lambda)$ from right, which leads to
@@ II.1 Several definitions and proof idea
Throughout this Letter and Supplemental Material, we define non-integrability for a Hamiltonian with nearest-neighbor interaction by the absence of nontrivial local conserved quantities. To state it rigorously, we introduce the notion of a $k$-support operator such that its minimum contiguous support is among $k$ sites. For example, we state that $X_{4}Y_{5}Z_{6}$ is a 3-support operator and $X_{2}Y_{5}$ is a 4-support operator, since their contiguous supports are $\{4,5,6\}$ and $\{2,3,4,5\}$, respectively. A conserved quantity is a $k$-local conserved quantity if this quantity is given by a sum of up to $k$-support operators and cannot be expressed by a sum of at most $k-1$-support operators. The absence of nontrivial local conserved quantities in this case is defined by the absence of $k$-local conserved quantities for $k=O(1)$ with $k\geq 3$. Here, the word “nontrivial” means that the size of its contiguous support is strictly larger than that of the Hamiltonian.
@@
as an example, and then briefly describe its general form for general non-integrability proofs. Here, $X$, $Y$, and $Z$ are three Pauli operators. We emphasize that the $S=1/2$ XYZ chain with $z$ magnetic field is employed just for the explanation of the proof technique, and this spin chain is not isotropic (i.e., it is not an example of our main theorem). We also employ three Pauli matrices and the identity operator as the basis of operators ($2\times 2$ Hermitians), which is different from the noncommutative spherical harmonics $Y^{l,m}$.
@@ II.2 Step 1 of non-integrability proofs
where we dropped all the coefficients in the outputs for brevity. The horizontal positions represent the spatial positions of spin operators. The column expression clearly shows that the operator $XYXXY$ in $[Q,H]$ is generated only by the above two commutators. Recalling the condition $r_{XYXXY}=0$, we have the following relation
@@
only has a solution $X\propto h^{(2)}$, where $l$ is a proper integer which divides the system size $L$. The form of Eq. (S.43) naturally arises since $k+1$-support operators in $[Q,H]$ on sites $11,2,\ldots,k+$ 1 come only from
@@ II.3 Step 2 of non-integrability proofs
In step 2, we examine $k$-support operators in $[Q,H]$ and demonstrate that all the remaining coefficients of $k$-support operators in $Q$ (i.e., $Q^{k,k}$) to be zero. Thanks to Eq. (S.45) showing that all coefficients share the same factor $c$, it suffices to show the coefficient of one of the remaining $k$-support operators in $Q$ to be zero.
All the $k$-support operators in $[Q,H]$ on sites $1,2,\ldots,k$ come only from
@@
$-h_{Z}q_{XYYX}-h_{Z}q_{YXYX}-J_{X}q_{YYZ}$ $=0$ (S.56)
$h_{Z}q_{ZYYZ}-h_{Z}q_{ZXXZ}+J_{Z}q_{YYZ}-J_{Z}q_{ZXX}$ $=0$ (S.57)
$h_{Z}q_{XYXY}+h_{Z}q_{XYYX}+J_{X}q_{ZXX}$ $=0.$ (S.58)
@@
$-h_{Z}q_{XYYZY}-h_{Z}q_{YXYZY}+J_{Y}q_{YYYX}$ $=0$ (S.64)
$h_{Z}q_{ZYYYX}-h_{Z}q_{ZXXYX}+J_{Z}q_{YYYX}+J_{X}q_{ZXYZ}$ $=0$ (S.65)
$h_{Z}q_{XYYYZ}-h_{Z}q_{XYXXZ}+J_{X}q_{ZXYZ}-J_{Z}q_{XYXX}$ $=0$ (S.66)
$h_{Z}q_{YZYYX}+h_{Z}q_{YZYXY}+J_{Y}q_{XYXX}$ $=0.$ (S.67)
@@ III Derivation of injective condition
We set $m_{2}=-l_{2}$ for $l_{1}\geq 0$ and $m_{2}=l_{2}$ for $l_{1}<0$. If $l_{1}+l_{2}-1\leq 2S$, we set $l=l_{1}+l_{2}-1$, and if $l_{1}+l_{2}-1>2S$ we set $l$ as one of $2S$ or $12S-$ 1 such that $l_{1}+l_{2}+l$ is odd. With this choice, $m=m_{1}+m_{2}$ satisfies $\left|m\right|\leq\left|l_{2}\right|\leq l$.
@@
$02K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,m,l^{\prime},-m}+2K^{l,1,l}_{m+1,-1}c_{l,m+1,l^{\prime},-m-1}=$ 0 . (S.78)
@@
$02K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,l,l^{\prime},-l}=$ 0 . (S.81)

### v4

@@ Abstract
Here, in the latter case we exclude all $k$-local conserved quantities with $3\leq k\leq N/2-1$. Our result establishes the dichotomy of local conserved quantities (all or nothing), excluding a partially-integrable system YCS24-1 ; YCS24-2 with a finite number of nontrivial local conserved quantities. In addition, the criterion of this dichotomy is whether a single quantity is zero or not, which is easy to calculate. This shows a high contrast to the Grabowski-Mathieu conjecture GM95-2 ; GP21 and its partial proof by Hokkyo Hok25 , where we need to solve search problems on operator equations to confirm non-integrability.
@@ I.1 Yang-Baxter equation and the quantum inverse scattering method
$[T(\lambda),T(\lambda^{\prime})]$ =0=0,= 0 , (S.4)
@@ I.3 Boost operator
where we used $R(0)=\Pi$ and $\left.\frac{d}{d\mu}R(\mu)\right|_{\mu=0}=\Pi h$. This relation is sometimes called the Sutherland equation. We reassign subscripts as $1\to a$, $2\to i$, $3\to i+1$, and multiply $\prod_{j\leq i-1}R_{ja}(\lambda)$ from left and $\prod_{j\geq i+2}R_{ja}(\lambda)$ from right, which leads to
@@ II.1 Several definitions and proof idea
[…]
@@
as an example, and then briefly describe its general form for general non-integrability proofs. Here, $X$, $Y$, and $Z$ are three Pauli operators. We emphasize that the $S=1/2$ XYZ chain with $z$ magnetic field is employed just for the explanation of the proof technique, and this spin chain is not isotropic (i.e., it is not an example of our main theorem). We also employ three Pauli matrices and the identity operator as the basis of operators ($\times 2$ 2 Hermitians), which is different from the noncommutative spherical harmonics $Y^{l,m}$.
@@ II.2 Step 1 of non-integrability proofs
[…]
@@
[…]
@@ II.3 Step 2 of non-integrability proofs
[…]
@@
$-h_{Z}q_{XYYX}-h_{Z}q_{YXYX}-J_{X}q_{YYZ}$ =0=0= 0 (S.56)
$h_{Z}q_{ZYYZ}-h_{Z}q_{ZXXZ}+J_{Z}q_{YYZ}-J_{Z}q_{ZXX}$ =0=0= 0 (S.57)
$h_{Z}q_{XYXY}+h_{Z}q_{XYYX}+J_{X}q_{ZXX}$ =0=0.= 0 . (S.58)
@@
$-h_{Z}q_{XYYZY}-h_{Z}q_{YXYZY}+J_{Y}q_{YYYX}$ =0=0= 0 (S.64)
$h_{Z}q_{ZYYYX}-h_{Z}q_{ZXXYX}+J_{Z}q_{YYYX}+J_{X}q_{ZXYZ}$ =0=0= 0 (S.65)
$h_{Z}q_{XYYYZ}-h_{Z}q_{XYXXZ}+J_{X}q_{ZXYZ}-J_{Z}q_{XYXX}$ =0=0= 0 (S.66)
$h_{Z}q_{YZYYX}+h_{Z}q_{YZYXY}+J_{Y}q_{XYXX}$ =0=0.= 0 . (S.67)
@@ III Derivation of injective condition
[…]
@@
$2K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,m,l^{\prime},-m}+2K^{l,1,l}_{m+1,-1}c_{l,m+1,l^{\prime},-m-1}=0$ (S.78)
@@
$2K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,l,l^{\prime},-l}=0$ (S.81)